The Data of Macroeconomics-Work It Out Question 2 Consider an economy that produces and consumes bread and automobiles. The table includes data for two different years. Good 20102010 2015 2015 QuantityriceQuantityPrice Automobiles| 70 $44000110 65000 Bread 530000$10 Round answers to to places afier the decimal where necessary a. Using 2010 as the base year, compute nominal GDP, real GDP, the implicit price deflator for GDP, and the CPI for each year 2010 Nominal GDP 2015 Nominal GDPS 2010 Real GDP S 2015 Real GDP 2010 GDP Deflator 2015 GDP Deflator 2010 CPI 2015 CPI

Need help, please show work
Answer #1

1) nominal gdp 2010 = quantity of automobiles * price of automobiles + quantity of bread * price of bread = 70 * $44,000 + 530,000 * $10 = $3,080,000 + $5,300,000 = $8,380,000

2) nominal gdp 2015 = quantity of automobiles * price of automobiles + quantity of bread * price of bread = 110 * $65,000 + 430,000 * $19 = $7,150,000 + $8,170,000 = $15,320,000

3) real gdp 2010 = quantity of automobiles * price of automobiles + quantity of bread * price of bread = 70 * $44,000 + 530,000 * $10 = $3,080,000 + $5,300,000 = $8,380,000

4) real gdp 2010 = quantity of automobiles * price of automobiles + quantity of bread * price of bread = 110 * $44,000 + 430,000 * $10 = $4,840,000 + $4,300,000 = $9,140,000